The average travel time from your residence to your school is 35 minutes with a standard deviation of 10 minutes. If you want to be 99% certain that you will not be late for first 8:00 am class, what is the latest time you should leave home?


1
Expert's answer
2021-05-31T15:18:13-0400

"P(Z<z)=0.99."

"z=2.326."

"\frac{x-35}{10}=2.326."

"x=2.325*10+35=58.25."

The latest time you should leave home: "8:00-58.25=7:01.75\approx7:02."